OpenStudy (anonymous):

Help please Factor: 24x2 - 23x - 12 A. (8x + 3)(3x - 4) B. (4x - 3)(6x + 4) C. (2x - 2)(12x + 6)

OpenStudy (anonymous):

Has this sir heard of synthetic division?

OpenStudy (anonymous):

No, this sir has not.

OpenStudy (anonymous):

what is the answer of 24*12

OpenStudy (anonymous):

288

OpenStudy (anonymous):

ok wait than

OpenStudy (anonymous):

STEPS TO BEING A SIR. STEP ONE: READ THY HANDBOOK http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

its simple 24x^2-32x+9x-12

OpenStudy (anonymous):

Yeah but where did you get the -32x+9x-12?

OpenStudy (anonymous):

24*12=288 so 32*9=288

OpenStudy (anonymous):

so we need -23 in centre so -32+9 gives us -23

OpenStudy (anonymous):

Oh okay. Now what?

OpenStudy (anonymous):

ok just tick the option A its right

OpenStudy (anonymous):

Yeah, but how'd you get it?

OpenStudy (anonymous):

24x^2-32x+9x-12=0 8x(3x-4)+3(3x-4)=0 (8x+3)(3x-4)

OpenStudy (anonymous):

understood now feel like sir

OpenStudy (anonymous):

One min, I'm working it out.

OpenStudy (anonymous):

How did you get this: 8x(3x-4)+3(3x-4)=0

OpenStudy (anonymous):

WAIT WAIT. I got it.

OpenStudy (anonymous):

by taking 8 common from 24 and 32 .....u know 8*3=24and 8*4=32

OpenStudy (anonymous):

Okay, I get it now. Thanks so much, man!